Contribution Title

Name of Author^{1} and Name of Author^{2}

1 Name and Address of your Institute name@email.address 2 Name and Address of your Institute name@email.address

Your text goes here. Separate text sections with the standard L^{A}TEX sectioning commands.

1 Section Heading

Your text goes here. Use the L^{A}TEX automatism for your citations [1].

1.1 Subsection Heading Your text goes here.

a × b = c  (1)

Subsubsection Heading Your text goes here. Use the L^{A}TEX automatism for cross-references as well as for your citations, see Sect. 1. Paragraph Heading Your text goes here. Subparagraph Heading. Your text goes here.

Table 1. Please write your table caption here

first second third number number number number number number

Fig. 1. Please write your figure caption here

Theorem 1. Theorem text^{3} goes here.

Lemma 1. Lemma text goes here.

References

1. Kipphan H (2000) Handbook of printmedia. Springer, Berlin Heidelberg New York

2. Brandt J, Hein W (2001) Polymer materials in joint surgery. In: Grellmann W, Seidler S (eds) Deformation and fracture behavior of polymers. Engineering materials. Springer, Berlin Heidelberg New York

3. Che M, Grellmann W, Seidler S (1997) Appl Polym Sci 64:1079–1090

4. Ross DW (1977) Lysosomes and storage diseases. MA Thesis, Columbia University, New York

3 Footnote

Index

paragraph, 1